OpenStudy (anonymous):

Which of the following are solutions to (2tanx)/(1- tan^2x) = sqrt(3) more than one avaliable a. x = (pi/3) + nPI b. x = (2pi/3) + nPI c x = (5pi/6) + npi d. x = (pi/6) + nPI

OpenStudy (anonymous):

d ption correct

OpenStudy (anonymous):

because tanx =tany then x=pi n+y

OpenStudy (anonymous):

and?

OpenStudy (anonymous):

you can notice that we are given the double angle formula for tan 2x so we can say that \[\tan 2x = \sqrt{3}\] we know that tan is only positive in the first and third quadrant on solving this \[\tan^{-1} \sqrt{3} = \frac{ \Pi }{ 3 }\] then the solution will be \[\frac{ \Pi }{ 6 }\]and its nPi multiplies so it will be d.

OpenStudy (anonymous):

right dear.......

OpenStudy (anonymous):

so is it just d ?

OpenStudy (anonymous):

yes doesn't it stated values of n ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

n pi what are the values of n ?

OpenStudy (anonymous):

i don't know?

OpenStudy (anonymous):

well it should be positive integers multiplies of 30 (pi/6)

OpenStudy (anonymous):

5pi/6 work?